### v1

# Uniform Feasibility For Smoothed Backup Control Barrier Functions

Abstract
We study feasibility guarantees for safety filters developed using Control Barrier Functions (CBFs) when a safe set is defined using the pointwise minimum of continuously differentiable functions, a construction that is common for the backup CBF method and typically nonsmooth. We replace the minimum by its log–sum–exp (soft-min) smoothing and show that, under a strict safety condition, the smooth function becomes a CBF (or extended CBF) for a range of the smoothing parameter. For compact safe sets, we derive an explicit lower bound on the smoothing parameter that makes the smooth function a CBF and hence renders the corresponding safety filter feasible. For unbounded sets, we introduce tail conditions under which the smooth function satisfies an extended CBF condition uniformly. Finally, we apply these results to backup CBFs. We show that safety of a compact (terminal) backup set under a backup controller, together with a condition ensuring safety of the backup trajectories on the relevant boundary of the safe set, is sufficient for feasibility for backup CBFs. These results provide a recipe for a priori feasibility guarantees for smooth inner approximations of nonsmooth safe sets without the need for additional online certification.
I Introduction
Safety is the main requirement for controlling modern autonomous systems. In safety-critical control, the objective is to provide a formal guarantee that, starting from a designated set, trajectories never leave this set during operation. Control Barrier Functions (CBFs) embed such guarantees as constraints in online optimization—often by penalizing the deviation from a desired control law, which leads to a design paradigm called safety filter [ames2019control]. The practicality of this framework ultimately hinges on the feasibility of the resulting optimization problem, which in turn requires a certificate that the constraint function is a valid CBF.
Converse safety theorems study conditions such that a safe set admits a CBF, therefore they can be used to address the certification problem when the safe set is represented by a continuously differentiable function. Converse results mostly rely on compactness assumption of the set, and span stability-like conditions [prajna2005ontheNecessity], separation certificate via Meyer functions [wisniewski2016converse], and robust variants [ratschan2018converse]. Alternatives avoid compactness assumption at the expense of other restrictive conditions such as global Lipschitz constant [liu2022converse] or bounded Lie derivatives [gurriet2018towards]. Recently, a relaxed formulation of CBF, named extended CBF (eCBF), has been proposed to remove these assumptions from the converse side of necessity (of the existence of an eCBF), while retaining safety guarantees on the direct side of sufficiency [mestres2024converse].
While converse safety theorems enable CBF certification, they presuppose the existence of a control law that keeps the system within the set, or equivalently, the set must be viable [^1] A set is viable if, from every point in it, the system can apply some admissible control that keeps the trajectory inside the set for all future time. under some control input in the sense of Aubin’s viability theory [aubin1991viability]. Ideally, one would find the largest viable set (viability kernel), but related computations are generally intractable in high dimensions due to the curse of dimensionality [mitchell2005aTime]. The backup control barrier function (BCBF) framework provides a scalable alternative [gurriet2018online, chen2021backup, wabersich2021predictive, breeden2022predictive]. Here, a backup controller is used to propagate states forward into a smaller viable “backup set,” and the set of all states that reach this backup set (at a terminal time) without leaving the safe region is implicitly defined via a minimum over time-indexed constraints, which leads to a nonsmooth. Since identifying a suitably small viable backup set can be achieved through convex relaxations or semidefinite programming, e.g., via LMIs [korda2013inner], this approach avoids direct computation of the viability kernel. The paradigm has been extended to data-driven settings via Koopman operators [folkestad2020data], to reduced-order models [molnar2023safety], to robust formulations [van2024disturbance], and to the multiple backup controllers setting [janwani2024learning]. However, feasibility guarantees in these BCBF-based methods often rely on ad hoc assumptions—such as presuming the existence of suitable class-$\mathcal{K}$ bounds without constructive certificates [gurriet2018online, chen2021backup], or imposing global compactness [gurriet2020scalable], or bounded Lie derivatives [rivera2024forward]—which limits systematic deployment.
Contributions. We bridge BCBFs with converse safety certification to find feasibility guarantees for CBF-based safety filters. We do this by smoothing the nonsmooth pointwise minimum via the soft-min (log–sum–exp) operator and provide a range of parameters such that converse safety theorems in [mestres2024converse] can be employed for CBF/eCBF certification. In particular: (i) For compact safe sets, uniform continuity can be invoked to provide CBF conditions. (ii) For noncompact sets, we introduce explicit tail conditions under which the soft-min uniformly satisfies an eCBF condition. (iii) In the BCBF setting, we show that safety of a compact backup set under a backup controller, combined with a condition enforcing safety along the relevant boundary of the safe set, is sufficient to certify the nonsmooth backup CBF as a CBF. This relaxes the global compactness assumption in [gurriet2020scalable]. While the smoothing idea also appeared in [rabiee2025soft] for BCBFs, the approach of [rabiee2025soft] requires computing an inner feasible subset online via an optimization problem for given design parameters. In contrast, our results provide an a priori and uniform feasibility guarantee that eliminates the need for additional online certification.
Organization. Section II reviews notation, (e)CBFs, converse safety, and the BCBF setup. Section III presents the soft-min construction and uniform thresholds for compact and noncompact cases. Section IV implements these results for BCBFs. Section V concludes the paper.
II Background
II-A Notation
We write $\mathbb{I}_{N}\triangleq\{1,\cdots,N\}$ for an integer $N\geq 1$. The set of continuously differentiable functions is denoted by $C^{1}$. For $h:\mathbb{R}^{n}\to\mathbb{R}$ and $h\in C^{1}$, its gradient is $\nabla h:\mathbb{R}^{n}\to\mathbb{R}^{n}$, and the transpose of its gradient is denoted as $\nabla^{\top}h$. A value $y\in\mathbb{R}$ is a regular value of $h$ if $\nabla h(x)\neq 0$ for every $x$ with $h(x)=y$. For a vector field $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$, $L_{f}h(x)\triangleq\nabla^{\top}\!h(x)f(x)$ denotes the Lie derivative of $h$ with respect to (w.r.t.) $f$. For a set $\mathcal{S}\subseteq\mathbb{R}^{n}$, $\partial\mathcal{S}$ and ${\rm Int}(\mathcal{S})$ denote its boundary and interior. A continuous function $\alpha:[0,a)\to\mathbb{R}_{\geq 0}$ belongs to class-$\mathcal{K}$ if it is strictly increasing and $\alpha(0)=0$. If $a=\infty$ and $\alpha(r)\to\infty$ as $r\to\infty$, then $\alpha\in\mathcal{K}_{\infty}$. A continuous function $\gamma:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ belongs to class-$\mathcal{K}_{\infty}\mathcal{K}$ if $\gamma(\cdot,s)\in\mathcal{K}_{\infty}$ for all $s\geq 0$ and $\gamma(r,\cdot)\in\mathcal{K}$ for all $r\geq 0$.
II-B Control Barrier Functions and Converse Safety Theorems
Consider a system
$\dot{x}=f(x,u),\quad x(0)=x_{0},$ (1)
with $x\in\mathbb{R}^{n}$ and $u\in\mathbb{U}\subseteq\mathbb{R}^{m}$, where $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$ is locally Lipschitz. We aim at safety, that the closed-loop system with a state-feedback controller $u=k(x)$ stays in a predefined set $\mathcal{S}$. We will consider sets defined as
$\begin{split}\mathcal{S}\triangleq\{x\in\mathbb{R}^{n}~|~h(x)\geq 0\},\quad h:\mathbb{R}^{n}\to\mathbb{R},~~h\in C^{1}.\end{split}$ (2)
Control Barrier Functions (CBFs) can be used to construct controllers with a safety guarantee.
Definition 1 (CBF, [ames2019control]).
A function $h\in C^{1}$ is a CBF for the system (1) if there exists a function $\alpha\in\mathcal{K}$ satisfying:
$\sup_{u\in\mathbb{U}}\Big[\underbrace{\nabla^{\top}\!h(x)f(x,u)}_{L_{f}h(x,u)}\Big]\geq-\alpha(h(x)),\quad\forall x\in\mathcal{S},$ (3)
where $L_{f}h$ denotes the Lie derivative of $h$ w.r.t. $f$.
We will utilize a less restrictive version of CBF called extended CBF to obtain more general results.
Definition 2 (Extended CBF, [mestres2024converse]).
A function $h\in C^{1}$ is an extended CBF for (1) if there exists a function $\gamma\in\mathcal{K}_{\infty}\mathcal{K}$ with:
$\sup_{u\in\mathbb{U}}\left[L_{f}h(x,u)\right]\geq-\gamma\left(h(x),\|x\|\right),\quad\forall x\in\mathcal{S}.$ (4)
CBFs and eCBFs imply that any locally Lipschitz continuous controller satisfying (3) or (4) ensures safety, provided that 0 is a regular value of $h$ [ames2019control, mestres2024converse]. One such controller that safeguards a desired (locally Lipschitz continuous) controller $k_{\rm des}$ is the safety filter:
$k_{\rm SF}(x)=$ $~\underset{u\in\mathbb{U}}{\rm argmin}~\|u-k_{\rm des}(x)\|$ (5)
$~\text{s.t.}~L_{f}h(x,u)\geq-\alpha(h(x)),$
where $\alpha(h(x))$ can be replaced with $\gamma(h(x),\|x\|)$ for eCBF. While certification of $h$ as a CBF or an eCBF implies the feasibility of the safety constraint, the optimization problem (5) becomes a QP (with efficient real-time solvers available) if $f$ is affine-in-control and $\mathbb{U}$ can be represented with affine constraints. Converse safety theorems certify $h$ as a CBF or an eCBF.
Theorem 1 (Converse safety for compact sets, [ames2019control]).
Let $\mathcal{S}$ be a compact set defined with $h\in C^{1}$ as in (2) and assume 0 is a regular value of $h$. If there exists a controller $k:\mathbb{R}^{n}\to\mathbb{U}$ such that $F(x)=f(x,k(x))$ is safe w.r.t. $\mathcal{S}$, then $h$ is a CBF.
Theorem 2 (Converse safety for general case, [mestres2024converse]).
Let $\mathcal{S}$ be a (potentially unbounded) set defined with $h\in C^{1}$ as in (2) and assume 0 is a regular value of $h$. If there exists a controller $k:\mathbb{R}^{n}\to\mathbb{U}$ such that the closed loop system $F$ is safe w.r.t. $\mathcal{S}$, then $h$ is an eCBF.
Remark 1.
The benefit of eCBFs with a more general class of functions $\gamma\in\mathcal{K}_{\infty}\mathcal{K}$ over $\alpha\in\mathcal{K}$ make them more suitable for unbounded sets. For example, even when $L_{F}h(x)$ becomes arbitrarily negative as $\|x\|\to\infty$, it is still plausible to construct a lower bound $\gamma(\cdot,\|x\|)$ using $L_{F}h$ thanks to the second term. However, missing this term, a CBF candidate may fail to admit a function $\alpha\in\mathcal{K}$. In particular, if $L_{F}h(x)\to$-$\infty$ as $\|x\|\to\infty$ and $h(x)=c\geq 0$ with a constant $c>0$, then no $\alpha\in\mathcal{K}$ can satisfy the barrier function condition with $F$ because $\alpha(c)$ is constant, see Example IV.1 in [mestres2024converse].
II-C Backup Control Barrier Functions
While converse safety theorems give us strong theoretical guarantees to certify any $h$ defining $\mathcal{S}$ as a CBF or an eCBF, they essentially require the existence of a control law rendering $\mathcal{S}$ safe. A scalable solution is proposed in [gurriet2018online]. This idea, which will be called backup control barrier functions in this paper, relies on constructing a (larger) safe set using finite-time safe trajectories that evolve within a (smaller) safe set $\mathcal{S}_{\rm b}\subseteq\mathcal{S}$.
Consider $h_{\rm b}:\mathbb{R}^{n}\to\mathbb{R}$, $h_{\rm b}\in C^{1}$ defining a set $\mathcal{S}_{\rm b}=\{x~|~h_{\rm b}(x)\geq 0\}$, and the closed-loop system with a locally Lipschitz continuous backup controller $k_{\rm b}:\mathbb{R}^{n}\to\mathbb{U}$:
$\dot{x}=F_{\rm b}(x)\triangleq f(x,k_{\rm b}(x)),~~x(0)=x_{0}.$ (6)
Solution trajectories to (6), denoted as $\phi(x_{0},\tau)$, can be used over a finite horizon $\tau\in[0,T]$ with a $T\geq 0$ to construct
$\mathcal{S}_{T}\!=\!\big\{x\in\mathbb{R}^{n}~|~h_{T}(x)\geq 0\big\},$ (7)
where
$h_{T}(x)\triangleq\min\left\{\min_{\tau\in[0,T]}h(\phi(x,\tau)),h_{\rm b}(\phi(x,T))\right\}.$ (8)
The set $\mathcal{S}_{T}$ contains all points $x\in\mathcal{S}$ such that, starting from $x$ at $\tau=0$, the system reaches $\mathcal{S}_{\rm b}$ by $\tau=T$ without leaving $\mathcal{S}$ at any $\tau\in[0,T]$ under the backup controller. Therefore, if $\mathcal{S}_{\rm b}$ is a safe set, so is $\mathcal{S}_{T}$, see [gurriet2020scalable, molnar2023safety] for a more formal statement.
The set $\mathcal{S}_{T}$ can potentially be used in converse safety theorems to certify $h_{T}$ as a CBF or an eCBF, which then guarantees the feasibility of the safety constraint in the safety filter with $h_{T}$. However, Theorems 1-2 require a continuously differentiable constraint function, whereas $h_{T}$ is not due to the minimum. To solve this problem, we will use a smooth inner approximation of $\min$ functions in $h_{T}$ in the next section. Before that, we use discretization for the minimum in $h_{T}$ over a continuous time interval $[0,T]$. In particular, consider a discretization $\tau_{i}\in\{0,\Delta\tau,\cdots,T\}$ with a uniform interval $\Delta\tau>0$ and $i\in\mathbb{I}_{N}$ with $N=T/\Delta\tau+1$. Discretization leads to an approximation of ${\mathcal{S}}_{T}$ defined with
$\hat{\mathcal{B}}\triangleq\{x\in\mathbb{R}^{n}~|~\hat{b}(x)\geq 0\},\quad\hat{b}(x)\triangleq\min_{i\in\mathbb{I}_{N}}b_{i}(x),$ (9)
where
$b_{i}(x)\triangleq\begin{cases}h\left(\phi(x,\tau_{i})\right),&~{\rm if}~i\in\mathbb{I}_{N}\setminus\{N\},\\ h_{\rm b}\left(\phi(x,T)\right),&~{\rm if}~i=N.\end{cases}$ (10)
We can replace the safety constraint of (5) with:
$L_{f}b_{i}(x,u)\geq-\alpha_{i}\left(b_{i}(x)\right),\quad\forall i\in\{1,\cdots,N\}.$ (11)
with an $\alpha_{i}\in\mathcal{K}$ for CBF and a $\gamma_{i}\in\mathcal{K}_{\infty}\mathcal{K}$ for eCBF. While (11) is sufficient for safety [^2] Technically, discretization yields a relaxation of the original safe set. While it is possible to shift the safety guarantee to $\hat{\mathcal{B}}$ by tightening constraints against imperfections emerging from discretization [gurriet2020scalable], we skip these details here to keep the focus on the merits of the approach presented. [he2025predictive], its feasibility depends on whether $\hat{b}$ can be certified as a CBF or an eCBF.
Our main goal in this paper is to obtain feasibility guarantees for CBF-based controllers such as (5) when the safety constraint is constructed as smoothened inner approximation of pointwise minimum of subconstraint functions. In the next section we set up this problem for a general setting and provide solutions for two separate cases: first, when the set $\hat{\mathcal{B}}$ is compact (using Theorem 1); then for a general case (using Theorem 2). Finally, we will investigate how these conditions translate back to the BCBF framework in Section IV.
III Main Results
III-A Problem Formulation
Let a set $\hat{\mathcal{S}}$ be defined as
$\hat{\mathcal{S}}=\{x\in\mathbb{R}^{n}~|~\hat{h}(x)\geq 0\},\quad\hat{h}(x)\triangleq\min_{i\in\mathbb{I}_{N}}h_{i}(x),$ (12)
with functions $h_{i}:\mathbb{R}^{n}\to\mathbb{R}$ and $h_{i}\in C^{1}$ for all $i\in\mathbb{I}_{N}$, and an integer $N\geq 1$. The active and inactive set of indices at a point $x\in\hat{\mathcal{S}}$ are denoted with $\mathcal{A}(x)\triangleq\{i\in\mathbb{I}_{N}~|~h_{i}(x)=\hat{h}(x)\}$ and $\mathcal{P}(x)\triangleq\mathbb{I}_{N}\setminus\mathcal{A}(x)$. We assume that the Mangasarian-Fromovitz constraint qualification (MFCQ) holds for $\partial\hat{\mathcal{S}}$, i.e., there exists a $v\in\mathbb{R}^{n}$ such that
$\nabla^{\top}\!h_{i}(x)\,v>0,\quad\forall i\in\mathcal{A}(x),~~\forall x\in\partial\hat{\mathcal{S}}$. (13)
[…]
$T_{\hat{\mathcal{S}}}(x)=\{v\in\mathbb{R}^{n}~|~\nabla^{\top}\!h_{i}(x)\,v\geq 0,~~\forall i\in\mathcal{A}(x)\}.$ (14)
When $h_{i}\in C^{1}$, MFCQ is satisfied whenever all the active gradients $\nabla h_{i}$ do not point in exactly opposite directions, so a common direction can increase all active constraints simultaneously.
We are interested in a case where a locally Lipschitz continuous controller $k:\mathbb{R}^{n}\to\mathbb{U}$ endows the system
$\dot{x}=F(x)=f(x,k(x)),\quad x(0)=x_{0},$ (15)
with a safety guarantee w.r.t. $\hat{\mathcal{S}}$. Before introducing the problem setting, we formally define safety.
Definition 3 (Strict safety).
A system (15) is strictly safe w.r.t. a set $\hat{\mathcal{S}}$ if
$\forall x_{0}\in\hat{\mathcal{S}}\implies x(t)\in{\rm Int}(\hat{\mathcal{S}}),~\forall t>0.$ (16)
It is noted that the safety definition in Definition 3 is a stricter notion than the safety definition commonly used within the CBF framework [ames2019control]. In particular, inspired by the strict invariance in [aubin1991viability], we consider cases where any trajectory on the boundary cannot stay on the boundary, and it is strictly pushed inside. Defined on a stricter condition, Theorems 1-2 are still applicable for the safety definition in Definition 3.
According to the strict invariance result in Aubin [aubin1991viability, $\S 4.3$], $F(x)$ must belong to the relative interior of the contingent cone $T_{\hat{\mathcal{S}}}(x)$ for all $x\in\partial\hat{\mathcal{S}}$ to ensure strict safety. When $\hat{\mathcal{S}}$ is defined as in (12), and under MFCQ, the strict invariance condition becomes equivalent to [clarke1998nonsmooth]
$L_{F}h_{i}(x)>0,\quad\forall i\in\mathcal{A}(x),\ \forall x\in\partial\hat{\mathcal{S}}$. (17)
Soft-min: To employ converse safety theorems, we construct a smooth inner approximation of $\hat{\mathcal{S}}$ using the soft-min (or log-sum-exp minimum) function:
$\tilde{\mathcal{S}}_{\theta}\!=\!\{x\in\mathbb{R}^{n}~|~\tilde{h}_{\theta}(x)\!\geq\!0\},~\tilde{h}_{\theta}(x)\!\triangleq\!-\frac{1}{\theta}\log\left(\sum_{i=1}^{N}{\rm e}^{-\theta h_{i}(x)}\right),$ (18)
where $\theta>0$ is a parameter adjusting the transition between two extreme cases: $\theta\to 0$ yielding equal contribution from all $h_{i}$, and $\theta\to\infty$ returning to only active constraints. Since $h_{i}\in C^{1}$ for all $i\in\mathbb{I}_{N}$, we can find the gradient:
$\nabla\tilde{h}_{\theta}(x)=\sum_{i=1}^{N}w_{\theta,i}(x)\nabla h_{i}(x),~w_{\theta,i}(x)\triangleq\frac{{\rm e}^{-\theta h_{i}(x)}}{\sum_{j=1}^{N}{\rm e}^{-\theta h_{j}(x)}},$ (19)
where $w_{\theta,i}:\mathbb{R}^{n}\to(0,1)$ are weights of how much each gradient $\nabla h_{i}$ contributes to $\nabla\tilde{h}_{\theta}$. Note $\sum_{i\in\mathbb{I}_{N}}w_{\theta,i}(x)=1$ for all $x\in\hat{\mathcal{S}}$.
Problem: Given that (17) holds (implied by strict safety w.r.t. $\hat{\mathcal{S}}$ and MFCQ), our particular goal is to show that there exists a $\theta$ such that we can use Theorems 1-2 to certify $\tilde{h}_{\theta}$ in (18) as a CBF or an eCBF. More precisely, we look for a uniform $\theta^{*}>0$ such that the following holds
$\nabla^{\top}\!\tilde{h}_{\theta}(x)F(x)=L_{F}\tilde{h}_{\theta}(x)>0,\quad\forall x\in\partial\tilde{\mathcal{S}}_{\theta},~\forall\theta>\theta^{*}.$ (20)
The strict positiveness of the Lie derivative $L_{F}\tilde{h}_{\theta}$ in (20) implies that $\nabla\tilde{h}_{\theta}(x)\!\neq\!0$ for all $x\in\partial\tilde{\mathcal{S}}_{\theta}$; so 0 is a regular value of $\tilde{h}_{\theta}$, and $\partial\tilde{\mathcal{S}}_{\theta}$ is a $C^{1}$ hypersurface. Therefore, (20) becomes sufficient for Aubin’s strict invariance condition [aubin1991viability], and Theorems 1-2 apply.
We address the problem of obtaining (20) by considering two cases separately. First, when $\hat{\mathcal{S}}$ is compact, we establish the existence of a $\theta^{*}$ without requiring any additional assumptions. Subsequently, we examine the case in which $\hat{\mathcal{S}}$ is unbounded and show that a uniform $\theta^{*}$ can still be obtained, provided that additional conditions on the subconstraint functions $h_{i}$ and the closed-loop vector field $F$ are satisfied.
Remark 2.
While both our approach and that of [rabiee2025soft] employ the same type of smooth approximation (18) for the minimum operator (12), the nature of the resulting feasibility guarantees differs. In [rabiee2025soft], feasibility of the safety filter is only ensured on a subset $\Gamma\subseteq\tilde{\mathcal{S}}_{\theta}$, where $\Gamma$ is determined through an auxiliary optimization problem. In our work, we instead derive conditions under which feasibility holds on $\tilde{\mathcal{S}}_{\theta}$, and these conditions are obtained prior to implementation, without the need for additional online certification.
III-B Compact Sets
In this part, we will consider the case that $\hat{\mathcal{S}}$ is a compact set. We start by defining an $\varepsilon$-tube around $\partial\hat{\mathcal{S}}$ as
$T_{\varepsilon}=\{x\in\hat{\mathcal{S}}~|~0\leq\hat{h}(x)\leq\varepsilon\}.$ (21)
Note that, for $\varepsilon>0$, we have $T_{\varepsilon}\subseteq\hat{\mathcal{S}}$. A property of soft-min functions is a lower bound:
$\hat{h}(x)-\frac{\log N}{\theta}\leq\tilde{h}_{\theta}(x),\quad\forall x\in\hat{\mathcal{S}},~\forall\theta>0.$ (22)
The lower bound in (22) implies the existence of a $\theta$ such that the boundary of $\tilde{\mathcal{S}}_{\theta}$ is fully contained in the $\varepsilon$-tube around $\partial\hat{\mathcal{S}}$; to be precise, $\partial\tilde{\mathcal{S}}_{\theta}\subset T_{\varepsilon}\subseteq\hat{\mathcal{S}}$ for all $\theta\geq\tfrac{\log N}{\varepsilon}$. Therefore, given a $\theta\geq\tfrac{\log N}{\varepsilon}$, showing that $L_{F}\tilde{h}_{\theta}(x)>0$ holds for all $x\in T_{\varepsilon}$ is sufficient for (20). Furthermore, since $\hat{\mathcal{S}}$ is compact, so are $T_{\varepsilon}$ and $\partial\tilde{\mathcal{S}}_{\theta}$ for an $\varepsilon>0$ and a $\theta>0$.
We can decompose the Lie derivative in (20):
$L_{F}\tilde{h}_{\theta}(x)\!=\!\sum_{\!\!\!i\in\mathcal{A}(x)\!\!}w_{\theta,i}(x)L_{F}h_{i}(x)\!+\!\sum_{\!\!\!j\in\mathcal{P}(x)\!\!}w_{\theta,j}(x)L_{F}h_{j}(x),$ (23)
where $w_{\theta,(\cdot)}$ are the weights given as in (19), characterizing how much the Lie derivative of each subconstraint contributes to the total Lie derivative $L_{F}\tilde{h}_{\theta}$. Notice that the summation is separated between active and inactive constraints, cf. (19), as this will be useful to check $L_{F}\tilde{h}_{\theta}(x)>0$ for all $x\in T_{\varepsilon}$. In particular, in order for the strict positiveness to hold, we need certain bounds on the terms in (23), which are presented in the next lemma. Before presenting this, consider the following notation defining the gap between a subconstraint function $h_{j}$ and the pointwise minimum: $\delta_{j}(x)\triangleq h_{j}(x)-\hat{h}(x)$. Notice that by definition we have $\delta_{i}(x)=0$ for all active constraints $i\in\mathcal{A}(x)$, whereas $\delta_{j}(x)>0$ for inactive ones $j\in\mathcal{P}(x)$.
Lemma 1.
Let $\hat{\mathcal{S}}$ defined as in (12) be compact with $h_{i}\in C^{1},~\forall i\in\mathbb{I}_{N}$ with $N\geq 1$, and assume MFCQ holds for $\partial\hat{\mathcal{S}}$. If the system (15) is strictly safe w.r.t. $\hat{\mathcal{S}}$, then there exists a $T_{\varepsilon}\subseteq\hat{\mathcal{S}}$ with an $\varepsilon>0$ such that following conditions hold $\forall x\in T_{\varepsilon}$ with bounds $M,r,d>0$
$\max_{i\in\mathbb{I}_{N}}\left|L_{F}h_{i}(x)\right|$ $\leq M,$ (24)
$\min_{i\in\mathcal{A}(x)}L_{F}h_{i}(x)$ $\geq r>0,$ (25)
$\min_{j\in\mathcal{P}(x)}\delta_{j}(x)$ $\geq d>0,$ (26)
Proof.
1) Proof of (24). Notice that $T_{\varepsilon}\subseteq\hat{\mathcal{S}}$ is compact as $\hat{\mathcal{S}}$ is compact. As $h_{i}\in C^{1}$ and $F$ is continuous, the map $x\mapsto L_{F}h_{i}(x)$ is continuous for each $i\in\mathbb{I}_{N}$, and therefore bounded on the compact set $T_{\varepsilon}$. Then (24) follows with
$M\triangleq\max_{i\in\mathbb{I}_{N}}\;\max_{x\in T_{\varepsilon}}|L_{F}h_{i}(x)|.$ (27)
2) Proof of (25). Define a set
$\Gamma\triangleq\{\,(x,i)~|~x\in\partial\hat{\mathcal{S}},\ i\in\mathcal{A}(x)\}.$ (28)
The strict safety condition and MFCQ yields the strict positiveness of the Lie derivative (17), i.e., $L_{F}h_{i}(x)>0,\forall(x,i)\in\Gamma$. Since $\partial\hat{\mathcal{S}}$ is compact and $\mathcal{A}(x)\subseteq\mathbb{I}_{N}$ is finite for every $x\in\partial\hat{\mathcal{S}}$, $\Gamma$ is compact. Because each $L_{F}h_{i}$ is continuous, the map $(x,i)\mapsto L_{F}h_{i}(x)$ is continuous on $\Gamma$ and therefore attains a strictly positive minimum:
$r_{0}\triangleq\min_{(x,i)\in\Gamma}L_{F}h_{i}(x)>0,$ (29)
By the continuity of the map $L_{F}h_{i}$, we may choose an $\varepsilon>0$ sufficiently small such that
$L_{F}h_{i}(x)\geq r_{0}/2>0,\quad i\in\mathcal{A}(x),~x\in T_{\varepsilon}.$ (30)
Then (25) follows with $r=r_{0}/2$.
3) Proof of (26). Define a set
$\Lambda\triangleq\{\,(x,j)~|~x\in\partial\hat{\mathcal{S}},~j\in\mathcal{P}(x)\,\}.$ (31)
Similar to $\Gamma$ as in the previous case, $\Lambda$ is a compact set. The map $(x,j)\mapsto\delta_{j}(x)$ is continuous and strictly positive on $\Lambda$ (by definition), so it attains a minimum:
$d_{0}\triangleq\min_{(x,j)\in\Lambda}\delta_{j}(x)>0.$ (32)
By the continuity of the map $\delta_{j}$, we may choose an $\varepsilon>0$ sufficiently small such that
$\delta_{j}(x)\geq d_{0}/2>0,\quad j\in\mathcal{P}(x),~x\in T_{\varepsilon}.$ (33)
Then (26) follows with $d=d_{0}/2$. ∎
Using the bounds from Lemma 1, we can now specify a range of $\theta$ such that $L_{F}\tilde{h}_{\theta}(x)>0$ holds for all $x\in T_{\varepsilon}$, which implies our goal of (20).
Theorem 3 ($\tilde{h}_{\theta}$ is a CBF for compact sets).
Let the assumptions of Lemma 1 hold with $\varepsilon,M,r,d>0$. Then, $\tilde{h}_{\theta}$ is a CBF for (1) for all $\theta>\theta^{*}$ where
$\theta^{*}=\max\left\{\frac{\log N}{\varepsilon},\frac{1}{d}\log{\left(\frac{N(r+M)}{r}\right)}\right\},$ (34)
where $N$ is the number of subconstraint functions $h_{i}$.
Proof.
Recall the decomposition of the Lie derivative $L_{F}\tilde{h}_{\theta}$ between active and inactive constraints as in (23). Using the bounds $M,r,d>0$ from Lemma 1, one can obtain
$L_{F}\tilde{h}_{\theta}(x)\geq r\sum_{i\in\mathcal{A}(x)}w_{\theta,i}(x)-M\sum_{j\in\mathcal{P}(x)}w_{\theta,j}(x),~~\forall x\in T_{\varepsilon}.$ (35)
The gap bound $\delta_{j}(x)\geq d$ implies (since $\theta>0$)
$h_{j}(x)\geq\hat{h}(x)+d\implies{\rm e}^{-\theta h_{j}(x)}\leq{\rm e}^{-\theta(\hat{h}(x)+d)}.$ (36)
Using (36) and (19) one gets
$\sum_{j\in\mathcal{P}(x)}w_{\theta,j}(x)\leq N{\rm e}^{-\theta d}.$ (37)
Since $\sum_{i\in\mathcal{A}(x)}w_{\theta,i}(x)=1-\sum_{j\in\mathcal{P}(x)}w_{\theta,j}(x)$, (35) and (37) yield
$L_{F}\tilde{h}_{\theta}(x)\geq r-(r+M)Ne^{-\theta d},~~\forall x\in T_{\varepsilon}.$ (38)
To ensure strict positiveness, we require
$r-(r+M)N{\rm e}^{-\theta d}>0\iff{\rm e}^{-\theta d}<\frac{r}{N(r+M)},$ (39)
which holds whenever
$\theta>\theta^{*}_{\rm core}\triangleq\frac{1}{d}\log{\left(\frac{N(r+M)}{r}\right)}.$ (40)
Notice that $\theta^{*}_{\rm core}$ is the second term in (34).
The soft-min lower bound (22) implies that choosing $\theta\geq\tfrac{\log{N}}{\varepsilon}$ ensures $\partial\tilde{\mathcal{S}}_{\theta}\subset T_{\varepsilon}$. Notice $\tfrac{\log{N}}{\varepsilon}$ is the first term in (34). Thus, choosing $\theta>\theta^{*}$ with $\theta^{*}$ as in (34) satisfies (20).
Finally, the strict positiveness of the Lie derivative implies $\nabla\tilde{h}_{\theta}(x)\neq 0$ on $\partial\tilde{\mathcal{S}}_{\theta}$, and thus 0 is a regular value of $\tilde{h}_{\theta}$. Then $\tilde{\mathcal{S}}_{\theta}$ is strictly safe based on Aubin’s strict invariance condition [aubin1991viability]; and by Theorem  1, $\tilde{h}_{\theta}$ is a CBF for all $\theta>\theta^{*}$. ∎
III-C General Case
In this part we study a set of sufficient conditions implying similar results without relying on the compactness property. In particular, since $T_{\varepsilon}$ may be unbounded, the condition (17) is no longer sufficient to satisfy (20) as the uniform bounds (24)-(26) are not guaranteed. Therefore, we distinguish between the portion where the state remains within a bounded region and the portion where $\|x\|$ grows unbounded. We refer to the former as the core and the latter as the tail. This distinction is useful because the core behaves as a compact set, so the arguments and uniform bounds from the previous section already apply, whereas the tail requires additional growth conditions to ensure a uniform lower bound on $L_{F}\tilde{h}_{\theta}$.
To formalize this, we decompose $T_{\varepsilon}$ using a radius $R>0$:
$T_{\varepsilon}=(T_{\varepsilon}\cap B_{R})\cup(T_{\varepsilon}\cap B_{R}^{\rm c}),\quad B_{R}\triangleq\{x\in R^{n}~|~\|x\|\leq R\},$ (41)
where $B_{R}^{\rm c}$ denotes the complement of $B_{R}$. The set $T_{\varepsilon}\cap B_{R}$ forms the compact core, whereas $T_{\varepsilon}\cap B_{R}^{\rm c}$ forms the tail, which is unbounded for any $R>0$ (if $\hat{\mathcal{S}}$ is unbounded). Therefore, we introduce additional assumptions to limit the contribution of inactive terms in $L_{F}\tilde{h}_{\theta}$ for the tail.
Assumption 1.
There exist constants $\varepsilon,R>0$ such that following statements are true for every $x\in T_{\varepsilon}\cap B_{R}^{\rm c}$:
  1. (a)
    There exists a nondecreasing function $\eta:[R,\infty)\to\mathbb{R}_{>0}$ with $\eta(r)\to\infty$ as $r\to\infty$ that satisfies
    $\min_{j\in\mathcal{P}(x)}\delta_{j}(x)\geq\eta(\|x\|).$ (42)
  2. (b)
    There exist constants $C,p\geq 0$ satisfying
    $\max_{j\in\mathcal{P}(x)}|L_{F}h_{j}(x)|\leq C(1+\|x\|)^{p}.$ (43)
  3. (c)
    There exists a constant $r_{\infty}>0$ satisfying
    $\min_{i\in\mathcal{A}(x)}L_{F}h_{i}(x)\geq r_{\infty}.$ (44)
Assumption 1 essentially assures that the active part of $L_{F}\tilde{h}_{\theta}$ (the first term in (23)) overcomes the inactive part (the second term) on the tail part ($\forall x\in T_{\varepsilon}\cap B_{R}^{\rm c}$). In particular, Assumption 1 (a) implies that $w_{\theta,j}\to 0$ for all inactive constraints $j\in\mathcal{P}(x)$ as $\|x\|\to 0$ with an exponential rate. Therefore, it will dominate the polynomial growth of $L_{F}h_{j}$ in this portion (per Assumption 1 (b)). As a result, the inactive part in (23) goes to 0 as $\|x\|\to\infty$ or $\theta\to\infty$. At any fixed $x$, we can find a finite $\theta^{*}$ such that the inactive term in (23) falls within the strictly positive margins of the uniform bound of the active part (that is, $r_{\infty}$ as assumed by Assumption 1 (c)).
Remark 3.
Assumption 1 (a) holds if (i) there is a unique minimizer for $\hat{h}$ at the tail end; i.e., $\exists i^{*}\in\mathbb{I}_{N},\mathcal{A}(x)\equiv i^{*},\forall x\in T_{\varepsilon}\cap B_{R}^{\rm c}$, and (ii) the gap $\delta_{j}$ is radially coercive for inactive constraints; i.e., $\delta_{j}(x)\to\infty$ as $\|x\|\to\infty$ holds for all $j\in\mathcal{P}(x)$. Under these conditions the function
$\eta(r)\triangleq\min_{j\neq i^{*}}\min_{\|x\|=r}\Big(h_{j}(x)-h_{i^{*}}(x)\Big)$ (45)
is well-defined, and it satisfies requirements in Assumption 1 (a). Failing to satisfy either of these may result arbitrarily small gap $\delta_{j}\to 0$ as $\|x\|\to\infty$, which prevents the dominance of the active part in $L_{F}\tilde{h}_{\theta}$.
Theorem 4 ($\tilde{h}_{\theta}$ is an eCBF for general case).
Let $\hat{\mathcal{S}}$ be defined as in (12) with $h_{i}\in C^{1},\forall i\in\mathbb{I}_{N}$ with $N\geq 1$, assume MFCQ holds for $\partial\hat{\mathcal{S}}$, and suppose that Assumption 1 holds. If the system (15) is strictly safe w.r.t. $\hat{\mathcal{S}}$, then there exists a $\theta^{*}>0$ such that $\tilde{h}_{\theta}$ is an eCBF for (1) for all $\theta>\theta^{*}$.
Proof.
We analyze the core and tail regions separately.
Core region $(T_{\varepsilon}\cap B_{R})$. The proof of Theorem 3 yields a critical value $\theta^{*}_{\rm core}$ as in (40) with terms $r,M,d>0$ whose existence is guaranteed per Lemma 1 as $T_{\varepsilon}\cap B_{R}$ is compact for any given $R>0$.
Tail region $(T_{\varepsilon}\cap B_{R}^{\rm c})$. Recall the decomposition (23). Assumption 1 (b)-(c) give, for all $x\in T_{\varepsilon}\cap B_{R}^{\rm c}$,
$L_{F}\tilde{h}_{\theta}(x)\geq r_{\infty}\!\!\sum_{i\in\mathcal{A}(x)}w_{\theta,i}(x)-C(1+\|x\|)^{p}\!\!\!\sum_{j\in\mathcal{P}(x)}w_{\theta,j}(x).$ (46)
The radially coercive gap $\min_{j\in\mathcal{P}(x)}\delta_{j}(x)\geq\eta(\|x\|)$ by Assumption 1 (a) can be used in a similar way to (36)-(37), which yields
$\sum_{j\in\mathcal{P}(x)}w_{\theta,j}(x)\leq(N-1)e^{-\theta\,\eta(R)},\quad x\in T_{\varepsilon}\cap B_{R}^{\rm c}.$ (47)
Using $\sum_{i\in\mathcal{A}(x)}w_{\theta,i}(x)=1-\sum_{j\in\mathcal{P}(x)}w_{\theta,j}(x)$, we obtain for all $x\in T_{\varepsilon}\cap B_{R}^{\rm c}$,
$L_{F}\tilde{h}_{\theta}(x)\geq r_{\infty}-(N-1)\big(r_{\infty}+C(1+R)^{p}\big)e^{-\theta\,\eta(R)}.$ (48)
The strict condition below ensures strict positivity in the tail:
$r_{\infty}>(N-1)\big(r_{\infty}+C(1+R)^{p}\big)e^{-\theta\,\eta(R)}\ \Longleftrightarrow\ \theta>\theta^{*}_{\mathrm{tail}},$ (49)
where
$\theta^{*}_{\mathrm{tail}}\triangleq\frac{1}{\eta(R)}\log\!\Bigg(\frac{(N-1)\big(r_{\infty}+C(1+R)^{p}\big)}{r_{\infty}}\Bigg).$ (50)
Finally, per (22), $\theta\geq\tfrac{\log{N}}{\varepsilon}\implies\partial\tilde{\mathcal{S}}_{\theta}\subset T_{\varepsilon}$. Therefore, choosing $\theta>\theta^{*}$ satisfies (20), where
$\theta^{*}=\max\!\left\{\frac{\log N}{\varepsilon},\theta^{*}_{\mathrm{core}},\theta^{*}_{\mathrm{tail}}\right\}.$ (51)
The strict positiveness of the Lie derivative implies $\nabla\tilde{h}_{\theta}(x)\neq 0$ on $\partial\tilde{\mathcal{S}}_{\theta}$, and thus 0 is a regular value of $\tilde{h}_{\theta}$. Then $\tilde{\mathcal{S}}_{\theta}$ is strictly safe based on Aubin’s strict invariance condition [aubin1991viability]; and by Theorem 2, $\tilde{h}_{\theta}$ is an eCBF for all $\theta>\theta^{*}$.
IV Connection to the Backup CBF Framework
We now connect the general results of Section III to the backup CBF setting introduced in Section II-C. Let a (open-loop) system $f\in C^{1}$ as in (1), a set $\mathcal{S}$ and a backup set $\mathcal{S}_{\rm b}$ defined with functions $h,h_{\rm b}\in C^{1}$, respectively. Suppose that a backup controller $k_{\rm b}\in C^{1}$ yields a closed-loop vector field $F_{\rm b}(x)=f(x,k_{\rm b}(x))\in C^{1}$, where the flow is denoted as $\phi(x_{0},\tau)$ with an initial condition $x(0)=x_{0}\in\mathcal{S}$. Recall from Section II-C
$\hat{\mathcal{B}}=\{x\in\mathbb{R}^{n}~|~\hat{b}(x)\geq 0\},\quad\hat{b}(x)=\min_{i\in\mathbb{I}_{N}}b_{i}(x),$ (52)
where
$b_{i}(x)=\begin{cases}h(\phi(x,\tau_{i})),&~{\rm if}~i\in\mathbb{I}_{N}\setminus\{N\},\\ h_{\rm b}(\phi(x,T)),&~{\rm if}~i=N\end{cases}$ (53)
for a finite horizon $T>0$ and a sampling interval $\Delta\tau>0$ with $\tau_{i}\in\{0,\Delta\tau,\cdots,T\}$ for all $i\in\mathbb{I}_{N}$ with $N=T/\Delta\tau+1$. We use the soft-min inner-approximation with a $\theta>0$:
$\tilde{b}_{\theta}(x)\triangleq-\frac{1}{\theta}\log\Big(\sum_{i\in\mathbb{I}_{N}}e^{-\theta b_{i}(x)}\Big).$ (54)
Our goal is to find conditions certifying $\tilde{b}_{\theta}$ as a CBF.
We start with defining the backward-expanded set of the (terminal) backup set $\mathcal{S}_{\rm b}$:
$\mathcal{B}_{N}\triangleq\{x\in\mathbb{R}^{n}~|~b_{N}(x)\triangleq h_{\rm b}(\phi(x,T))\geq 0\}.$ (55)
Note that $\mathcal{B}_{N}$ consists of points $x$ such that, starting from $x$ at $\tau=0$, the flow $\phi(x,\tau)$ reaches in $\mathcal{S}_{\rm b}$ at $\tau=T$. Moreover, $\mathcal{B}_{N}\supseteq\hat{\mathcal{B}}$. What follows are useful properties of $\mathcal{B}_{N}$.
Lemma 2.
Let $F_{\rm b},h_{\rm b}\in C^{1}$, let $\mathcal{S}_{\rm b}$ be a compact set, let 0 be a regular value of $h_{\rm b}$, assume that the flow $\phi(x,\tau)$ of $F_{\rm b}$ is uniquely defined for all $|\tau|\leq T$ and all $x\in\mathbb{R}^{n}$. If $F_{\rm b}$ is strictly safe w.r.t. $\mathcal{S}_{\rm b}$, then it is also strictly safe w.r.t. $\mathcal{B}_{N}$. Moreover, $\mathcal{B}_{N}$ is compact.
Proof.
(i) Compactness.  Define the map $\Phi_{\tau}(x)=\phi(x,\tau)$. As the solution flow uniquely exists on $\tau\in[-T,T]$ for all $x\in\mathbb{R}^{n}$, $\Phi_{\tau}$ is injective, because no two different initial conditions can lead to a same point in its image (if $\Phi_{\tau}(x_{1})=\Phi_{\tau}(x_{2})$, then $x_{1}=x_{2}$). Thus, $\Phi_{\tau}:\mathbb{R}^{n}\mapsto\Phi_{\tau}(\mathbb{R}^{n})$ is bijective for all $|\tau|\leq T$, where $\Phi_{\tau}(\mathbb{R}^{n})$ is its image. Furthermore, since $F_{\rm b}\in C^{1}$, $\Phi_{\tau}$ is $C^{1}$ for all $|\tau|\leq T$. Thus, $\Phi_{\tau}$ is a $C^{1}$ diffeomorphism with a $C^{1}$ inverse $\Phi_{-\tau}$ for all $|\tau|\leq T$ [lee2003smooth]. This implies that $\mathcal{B}_{N}$ is the continuous image of the compact set $\mathcal{S}_{\rm b}$ under the inverse: $\mathcal{B}_{N}=\Phi_{-T}(\mathcal{S}_{\rm b})$, and therefore $\mathcal{B}_{N}$ is compact.
(ii) Safety. We use the semigroup property of the flow similar to [gurriet2020scalable]—but for strict safety. Take any $x\in\mathcal{B}_{N}$, then $\Phi_{T}(x)\in\mathcal{S}_{\rm b}$ by definition of $\mathcal{B}_{N}$. Since the backup system is strictly safe w.r.t. $\mathcal{S}_{\rm b}$, for every $s>0$ we have $\Phi_{T+s}(x)\in{\rm Int}(\mathcal{S}_{\rm b})$. Thus, for every $s>0$, $b_{N}(\Phi_{s}(x))=h_{\rm b}(\Phi_{T+s}(x))>0$, so $\Phi_{s}(x)\in{\rm Int}(\mathcal{B}_{N})$. This is exactly the safety definition in Definition 3. ∎
Safety of $\mathcal{B}_{N}$ creates a trivial case. In particular, if $\mathcal{B}_{N}\subset{\rm Int}(\mathcal{S})$, then backup trajectories $\phi$ cannot visit $\partial\mathcal{S}$ as they have to stay inside $\mathcal{B}_{N}$, and this implies $\hat{\mathcal{B}}=\mathcal{B}_{N}$. In this case, Lemma 2 provides us with all that is required to invoke Theorem 3; so $\hat{b}(x)=b_{N}(x)$ can be certified as a CBF (without using $h$).
In the more general case of $\hat{\mathcal{B}}\subset\mathcal{B}_{N}$, we need to pay close attention to the boundary points of $\hat{\mathcal{B}}$ where discretized ‘slice’ [^4] By ‘slice’ we mean evaluating $h$ at a fixed time slice of the backup trajectory: the constraint $b_{i}(x)=h(\phi(x,\tau_{i}))$ records the value of $h$ after flowing the system from $x$ for $\tau_{i}$ seconds of time. constraints of $b_{i}(x)=h(\phi(x,\tau_{i}))$ are active for some $i<N$. Observe that $\phi(x,\tau_{i})\in\partial\mathcal{S}$ at some $x\in\partial\hat{\mathcal{B}}$ if a slice becomes active $b_{i}(x)=\hat{b}(x)$ for some $i<N$. Furthermore, since $\mathcal{B}_{N}$ is strictly safe under the backup flow, all such evaluations necessarily lie in $\mathcal{B}_{N}$ as well. Consequently, every potential activation of some $i<N$ on the boundary of $\hat{\mathcal{B}}$ occurs on the backup-reachable boundary of $\mathcal{S}$:
$\mathcal{C}\triangleq\partial\mathcal{S}\cap\mathcal{B}_{N}.$ (56)
Assumption 2.
The following holds:
$L_{F_{\rm b}}h(y)>0,~~\forall y\in\mathcal{C}.$ (57)
Assumption 2 translates as the strict safety wherever $\partial\mathcal{S}$ becomes relevant for the backup trajectory. Together with the strict safety of $\mathcal{S}_{\rm b}$, this yields the strict safety of $\hat{\mathcal{B}}$, which is required to invoke Theorem 3. Note that the condition is required only on $\mathcal{C}$ because this set consists of the boundary points of the target safe set $\mathcal{S}$ and the backward-expanded set of $\mathcal{B}_{N}$, i.e., exactly the states at which the backup-based construction is applied. Any other boundary point of $\mathcal{S}$ that is outside $\mathcal{C}$ is not relevant for the formulation as $x\notin\mathcal{B}_{N}$ implies $\phi(x,T)\notin\mathcal{S}_{\rm b}$.
Corollary 1.
Let $F_{\rm b},h,h_{\rm b}\in C^{1}$, let $\mathcal{S}_{\rm b}$ be compact, 0 is a regular value of $h_{\rm b}$, and suppose the flow $\phi(x,\tau)$ of $F_{\rm b}$ exists for all $|\tau|\leq T$ for every initial condition $x\in\hat{\mathcal{B}}$. If $F_{\rm b}$ is strictly safe w.r.t. $\mathcal{S}_{\rm b}$ and Assumption 2 holds, then there exists a $\theta^{*}>0$ such that $\tilde{b}_{\theta}$ is a CBF for (1) for all $\theta\geq\theta^{*}$.
Proof.
We verify the four hypotheses of Theorem 3 (see Lemma 1) in order.
$\hat{\mathcal{B}}$ is compact
By Lemma 2, $\mathcal{B}_{N}$ is compact and strictly safe under $F_{\rm b}$. For any $i<N$, we have $b_{i}(x)=h(\phi(x,\tau_{i}))$, which is continuous, hence $\mathcal{B}_{i}=\{x~|~b_{i}(x)\geq 0\}$ is closed. Therefore $\hat{\mathcal{B}}=\bigcap_{i=1}^{N}B_{i}\subseteq B_{N}$ is the intersection of a compact set with finitely many closed sets, and is thus compact.
$b_{i}\in C^{1}$ for all $i\in\mathbb{I}_{N}$
Since $F_{\rm b},h,h_{b}\in C^{1}$ and the flow exists for all $|\tau|\leq T$, the flow $x\mapsto\phi(x,\tau)$ is $C^{1}$ for each fixed $\tau$. Thus, subconstraint functions $b_{i}(x)=h(\phi(x,\tau_{i}))$ for a $i<N$ and $b_{N}(x)=h_{\rm b}(\phi(x,T))$ are all $C^{1}$. Indeed
$\nabla b_{i}(x)=\begin{cases}D_{x}^{\top}\phi(x,\tau_{i})\,\nabla h(\phi(x,\tau_{i})),&i<N,\\ D_{x}^{\top}\phi(x,T)\,\nabla h_{\rm b}(\phi(x,T)),&i=N,\end{cases}$ (58)
where $D^{\top}_{x}\phi\!\in\!\mathbb{R}^{n\times n}$ is the transpose of the Jacobian matrix of $\phi(x,\cdot)$ w.r.t. $x$.
MFCQ holds on $\partial\hat{\mathcal{B}}$
For an arbitrary $x\in\partial\hat{\mathcal{B}}$, we consider two cases: $i\in\mathcal{A}(x)$ for a $i<N$, or $N\in\mathcal{A}(x)$. The former implies $\phi(x,\tau_{i})\in\partial\mathcal{S}$. In addition, since $\hat{\mathcal{B}}\subseteq\mathcal{B}_{N}$, and $\mathcal{B}_{N}$ is strictly safe, then $\phi(x,\tau_{i})\in\mathcal{B}_{N}$. Then Assumption 2 applies, and we have that $L_{F_{\rm b}}b_{i}(x)>0$. Thus, taking $v=F_{\rm b}(x)$ satisfies (13) for all $x\in\partial\hat{\mathcal{B}}$ when $i$-th constraint is active with $i<N$. If $N\in\mathcal{A}(x)$ for a $x\in\partial\hat{\mathcal{B}}$, then $\phi(x,T)\in\partial\mathcal{S}_{\rm b}$, and strict safety of $\mathcal{S}_{\rm b}$ yields $L_{F_{\rm b}}h_{\rm b}(\phi(x,T))>0$ as 0 is a regular value of $h_{\rm b}$. Similarly, $v=F_{b}(x)$ satisfies (13) for this case. Thus, MFCQ holds on $\partial\hat{\mathcal{B}}$ with $v=F_{\rm b}(x)$.
(The backup system is strictly safe w.r.t. $\hat{\mathcal{B}}$
Under MFCQ, Aubin’s strict invariance condition [aubin1991viability] reduces to the strict positiveness of the Lie derivative on boundary, cf. (17). We verify this inequality at an arbitrary $x\in\partial\hat{\mathcal{B}}$ for the same two cases as in the previous part: $i\in\mathcal{A}(x)$ for a $i<N$, or $N\in\mathcal{A}(x)$. Case $i<N$, as discussed in the previous part, Assumption 2 applies, and $L_{F_{\rm b}}b_{i}(x)>0$ for all $x\in\partial\hat{\mathcal{B}}$ when $i$-th constraint is active with $i<N$. Case $i=N$, means $\phi(x,T)\in\partial\mathcal{S}_{\rm b}$ and, since the backup closed loop is strictly safe w.r.t. $\mathcal{S}_{\rm b}$ and 0 is a regular value for $h_{\rm b}$, we have $L_{F_{\rm b}}h_{\rm b}(\phi(x,T))>0$. Thus (17) holds, and strict safety follows from Aubin’s strict invariance condition [aubin1991viability].
Having established (a)–(d), all hypotheses of Theorem 3 are satisfied, and this concludes the proof. ∎
Remark 4.
Corollary 1 relies on the compactness of the backup set $\mathcal{S}_{\rm b}$; and we can still certify $\tilde{b}_{\theta}$ as a CBF even in the case $h$ defines an unbounded $\mathcal{S}$ (provided that Assumption 2 holds). This relaxes the stronger (global) compactness assumptions found in the literature; e.g., [gurriet2020scalable]. If the backup set $\mathcal{S}_{\rm b}$ is not compact, then we can still use Theorem 4 to certify $h$ as an eCBF, which would imply the feasibility of the safety filter (5) all the same. We leave this as a topic for future work.
V Conclusion
We have established uniform feasibility guarantees for CBF-based controllers such as safety filters using converse safety theorems when the safe set is constructed using smoothened minima of $C^{1}$ functions (via soft-min). For compact safe sets, an explicit parameter is designated for certifying the soft-min function as a CBF, which implies the feasibility of the CBF-based safety constraint. For noncompact sets, a set of tail conditions was shown to yield an extended CBF (a more relaxed formulation of CBF) with the same feasibility results. Projecting theoretical findings onto the backup CBFs, we show that strict safety of a compact backup set under a backup controller, combined with a strict safety condition along the relevant boundary of the safe set, is sufficient for feasibility in nonsmooth backup CBF formulation.
Future work will focus on relaxing assumptions of compact backup set and strict positiveness on the backup-reachable boundary of the safe set. Another direction is to investigate regularity (local Lipshitzness) and robustness properties of the smoothed BCBF framework, including sensitivity to model mismatch and sampling/discretization effects.